OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). (2 points) y = - x - 6 y = - x - 14 y = 3x - 6 y = 3x - 14

OpenStudy (zale101):

What is the slope of the given equation 3x-y=-3

OpenStudy (anonymous):

-3?

OpenStudy (zale101):

No, 3x-y=-3 is the standard form, to convert this equation into the slope intercept form y=mx+b, you should solve for y by adding 3 to both sides and adding y to both sides. 3x+3=y

OpenStudy (anonymous):

so the answer would be y = 3x - 6

OpenStudy (zale101):

So the question is looking for a parallel line that passes (4,-2). Parallel lines will have the same slope of the equation y=3x+3. Use the point slope formula y-y1=m(x-x1) and plug in 3 for the slope and (4,-2) for the points x1.y1. You will have the function that is parallel to the given function y=3x+3